# POMATO: Marrying Pointmap Matching with Temporal Motions for Dynamic 3D Reconstruction

### v1

@@ 3.4 Motion Module for Video Pointmap Matching
3D Reconstruction.Within the window of a video sequence, our POMATO can leverage the advantage of the motion module and perform 3D reconstruction in a feed-forward manner, which skips redundant post-process operations like global alignment. As shown in the bottom part of the Fig. 4, the input video sequence is formulated as $\{({\bf I}^{t_{k}},{\bf I}^{t_{1}}),({\bf I}^{t_{k}},{\bf I}^{t_{2}}),...,({\bf I}^{t_{k}},{\bf I}^{t_{T}})\}$, where ${\bf I}^{t_{k}}$ is the keyframe. The basic idea is to align the pointmap of all reference frames to the coordinate system of the keyframe, and thus all the reference frames are fed to the Head2 while the keyframe is fed to the Head1. With the advantage of the motion module, the pointmap of each reference frame maintains a consistent scale under the same coordinate of the keyframe. Temporal consistency loss, as mentioned in Eq6, is similarly required for both involved heads.
@@
Alignment Method Optim. Onl. Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$ <$1.25\uparrow$
@@
Table 1: Video depth evaluation. We report scale-invariant depth and scale &\&& shift invariant depth accuracy on Sintel [4], Bonn [28], and KITTI [12] datasets. Methods requiring global alignment are marked “GA”, while “Optim.” and “Onl.” indicate optimization-based and online methods, respectively. The best and second best results in each category are bold and underlined, respectively.
@@ 4.1 Experimental Details
Training data. We train our network with a mixture of five datasets: PointOdyssey [52], Tartanair [43], ParallelDomain4D [38], DynamicReplica [19] and Carla (0.9.15) [9]. The specific number and the usage ratio of each dataset can be found in the supplementary materials. All of these datasets are synthetic and come with pixel-accurate ground truth depth, as well as camera intrinsics and extrinsics. They feature diverse dynamic scene types: indoor and outdoor. Among them, PointOdyssey and DynamicReplica have additional 2D trajectory annotations which can be used to construct pointmap matching ground truth following Eq. 2. Thus, all datasets contribute to geometry supervision for training pointmaps in Head1 and Head2 and we use only PointOdyssey, DynamicReplica, and TartanAir datasets to train our proposed pointmap matching head. The training of POMATO follows a two-stage process. In the first stage, we utilize pair-wise data to establish the fundamental capabilities of both geometry and matching. All parameters in the decoders and each DPT head are learnable. In the second stage, we introduce the temporal motion module and freeze parameters in the decoder. The same datasets are used in both stages.
Training and inference details. Our model architecture is based on the publicly available DUSt3R [50] model, utilizing the same backbone consisting of a ViT-Large encoder and a ViT-Base decoder. To maximize the benefits of MonST3R’s geometry estimation ability in dynamic scenes, we initialize the model weights using the publicly available MonST3R checkpoint. For the newly introduced pointmap matching head, we initialize the head weights from the Head2 weight of the MonST3R. We train our network for 10 epochs with a cosine schedule and the initial learning rate is set to 0.0001. The batch size for the first stage of pairwise image training is 16 on 4 A100 (40G) GPUs. When training with the temporal motion module, the batch size is set to 4 with a temporal window length of 12.
@@ 4.2 Video Depth Estimation
Following MonST3R [50] and CUT3R [41], we rescale all predictions from the same video to align them together by conducting two forms of alignment: per-sequence scale and shift alignment and per-sequence scale alignment. Thus, we can measure the per-frame depth quality and inter-frame depth consistency. We employ our proposed motion module for video depth estimation as described in Sec.3.4 and compare our method against several variants of DUSt3R, including DUSt3R [42], MAST3R [24], MonST3R [50], Spann3R [39], and CUT3R [41]. Some of these methods [42, 24, 50] rely on global alignment (GA), which assumes a static environment. While GA benefits multi-view consistency in static regions, it can potentially degrade the representation of dynamic elements and is computationally expensive. In contrast, POMATO can perform online video depth inference with the introduced temporal fusion module and achieve much faster inference speed. As shown in Tab. 1, our method demonstrates comparable performance to the global alignment (GA)-based MonST3R [50] on the Sintel [4] and BONN [28] datasets, while surpassing it on KITTI dataset. Besides, we consistently outperform the state-of-the-art online method, CUT3R [41], across various settings. These results underscore the effectiveness of our approach, specifically (1) the joint learning of geometry and pointmap matching, and (2) the temporal motion module.
@@
Method T-12 T-24 T-12 T-24 T-12 T-24 T-12 T-24
SpatialTracker [45] 21.24 23.33 22.34 22.87 32.94 32.81 25.51 26.34
DUSt3R [42] 20.27 21.01 30.77 28.64 10.91 6.39 20.65 18.68
MASt3R [24] 17.36 18.29 27.18 25.51 12.83 7.73 19.12 17.17
MonST3R [50] 28.50 29.29 29.82 28.42 18.23 10.81 25.52 22.84
POMATO 34.19 35.75 33.41 30.72 27.05 23.81 31.55 30.09
Table 2: 3D tracking evaluation. We report the APD metric to evaluate 3D point tracking on the PointOdyssey [52], ADT [29], and PStudio [18] datasets. T-12 and T-24 indicate tracking within the temporal length of 12 frames and 24 frames, respectively. POMATO achieves remarkable performance even compared with the specialized model SpatialTracker [45]. * indicates the camera intrinsic is required.
@@
For 3D point tracking task, we use the Aria Digital Twin (ADT) [29], and Panoptic Studio (PStudio) [18] benchmarks from the TAPVid-3D [23] dataset along with the validation set on the PointOdyssey [52] dataset. We report the Average Percent Deviation (APD) metric, which quantifies the average percentage of points within a threshold relative to the ground truth depth. The APD metric serves as a direct measure of the accuracy of the predicted tracking. We reformulate the datasets and project all the query points within a testing sequence to the first frame. We report tracking results on a length of both 12 and 24 frames. As shown in Tab.2, our POMATO achieves the best performance on both PointOdyssey and ADT datasets. Additionally, our method demonstrates superior generalization, as reflected by the highest average APD metric. It’s worth mentioning that SpatialTracker [45] is a state-of-the-art network tailored for 3D point tracking with ground truth camera intrinsic as additional input data. POMATO surpasses it on two datasets and improves the average APD metric by 23.7% and 14.2% for 12 frames and 24 frames, respectively. In contrast, DuST3R-based methods struggle with ambiguous matching representations, resulting in imprecise tracking performance in dynamic scenarios.
Video Depth Tracking
Temporal Length Sintel [4] Bonn [28] KITTI [12] ADT [29] PStudio [18] PointOdyssey [52]
@@
Pair-wise 0.548 46.2 0.087 94.0 0.113 89.5 31.35 25.20 33.21
6 frames 0.436 51.3 0.076 95.9 0.085 93.5 32.63 26.93 33.88
12 frames 0.416 53.6 0.075 96.1 0.086 93.3 33.41 27.05 34.19
Table 3: Ablation study on the temporal motion module. The introduction of the temporal motion module brings a significant improvement compared with the fundamental model trained on pairwise images. As the temporal window length enlarges from 6 frames to 12 frames, we obtain an overall consistent improvement for both video depth estimation and 3D point tracking tasks.
@@
POMATO 0.020 0.010 0.509 0.037 0.016 1.837
Table 4: Pose estimation. * indicates using an off-the-shelf optical flow model to get the motion mask and the pseudo 2D optical flow ground truth. Our method achieves an overall best performance and improves the RPE rot metric significantly.
Figure 7: Effectiveness of our motion mask estimation in 3D reconstruction. Without explicitly filtering out the motion area, both pose and geometry estimation will be degenerated.
@@ 4.4 Camera Pose Estimation
We conduct pose estimation experiments on Bonn  [28] and TUM [35] datasets which include dynamic moving objects, and compare our method with the DUSt3R-related works along with the recently proposed CUT3R[41]. Global alignment is utilized for the evaluation. The sampling stride is set to 5 on the Bonn dataset and 3 on the TUM dataset. We evaluate over 40 frames and report the results in Tab. 4. Three metrics are reported: Absolute Translation Error (ATE), Relative Translation Error (RPE trans), and Relative Rotation Error (RPE rot). Our method achieves the best overall performance. Notably, POMATO operates without relying on any auxiliary modules for dynamic object estimation, thereby avoiding potential domain gaps across modules and preventing error accumulation. In particular, POMATO significantly improves the RPE-rot metric, surpassing MonST3R by 55.4% and 13.3% on the TUM and Bonn datasets, respectively. We present a visualization to demonstrate the effectiveness of our dynamic mask estimation in 3D reconstruction, as mentioned in Sec.3.3. Without explicitly identifying dynamic regions, both camera pose and geometry estimation suffer from significant degeneration.
@@ 4.5 Ablation Study
Extensive ablation experiments are conducted on video depth estimation and 3D point tracking tasks to validate the effectiveness of learning temporal motions on video sequence input. We report three models trained with pairwise images, a shorter temporal window length of 6 frames, and the default temporal window length of 12 frames. As shown in Tab. 3, the introduction of the temporal motion module introduces significant improvements across all datasets, emphasizing the importance of consistency among different frames. When enlarging the temporal window length from 6 frames to 12 frames, the video depth estimation obtains a further improvement on both Sintel and Bonn datasets. For 3D point tracking, the enlarged temporal window length brings a consistent enhancement across all three testing datasets.
@@ 5 Discussion and Conclusion
We introduce POMATO, a unified framework designed for geometry estimation and motion understanding in dynamic scenes. Leveraging our proposed pointmap matching head, our method can effectively distinguish moving regions, thereby mitigating the interference caused by dynamic objects. The introduced temporal motion module enhances the learning of temporal motions across different frames, improving the scale consistency and boosting the performance in tasks where geometry and matching are critical, most notably, 3D point tracking. Moving forward, we aim to explore methods to scale up our training with more matching data and further improve the 3D reconstruction performance.
@@
${\bf D}^{j,i}=||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||>\alpha,$ (7)
where $\alpha$ is a dynamic threshold defined as $3\times\text{median}(||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||)$.

### v2

@@ 3.5 Downstream Temporal Tasks
3D Reconstruction. Assisted by the temporal motion module, redundant post-process operations such as global alignment can be omitted, allowing the reconstructed 3D point cloud to be obtained in a feed-forward manner. As shown in the bottom part of Fig.5, the keyframe is set to the last frame ${\bf I}^{t_{T}}$ within the temporal window of length $T$ and is fed to Head1 with a set output of $\{{\bf X}^{t_{T},t_{T}},{\bf X}^{t_{T},t_{T}},...,{\bf X}^{t_{T},t_{T}}\}$. All the reference frames are input to the Head2 so the target pointmaps $\{{\bf X}^{t_{1},t_{T}},{\bf X}^{t_{2},t_{T}},...,{\bf X}^{t_{T},t_{T}}\}$ are aligned under the coordinate system of the keyframe. The temporal consistency loss $\mathcal{L}_{\text{t}}$ is:
$\mathcal{L}_{\text{t}}=\frac{1}{T}\sum_{i=1}^{T}\left\|\frac{{{\bf X}^{t_{T},t_{T}}}}{z^{T}_{1}}-\frac{{\bar{{\bf X}}^{t_{T},t_{T}}}}{\bar{z}_{1}^{T}}\right\|+\left\|\frac{{{\bf X}^{t_{i},t_{T}}}}{z_{2}^{T}}-\frac{{\bar{{\bf X}}^{t_{i},t_{T}}}}{\bar{z}_{2}^{T}}\right\|$ (8)
We further freeze the parameters in Decoder1 and Decoder2 when training the temporal downstream tasks at the second stage. In our work, the temporal window length $T$ is set to 12. Additional explorations on the temporal length can be found in Sec.4.
@@
Alignment Method Optim. Onl. Abs Rel $\downarrow$ $\delta$$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$ $1.25\uparrow$
@@
Table 1: Video depth evaluation. We report scale-invariant depth and scale &\& shift invariant depth accuracy on Sintel [4], Bonn [30], and KITTI [13] datasets. Methods requiring global alignment are marked “GA”, while “Optim.” and “Onl.” indicate optimization-based and online methods, respectively. The best and second best results in each category are bold and underlined, respectively.
@@ 4.1 Experimental Details
Training data. We train our network with a mixture of five datasets: PointOdyssey [54], Tartanair [45], ParallelDomain4D [40], DynamicReplica [21] and Carla (0.9.15) [10]. The specific number and the usage ratio of each dataset can be found in the supplementary materials. All datasets include pixel-accurate ground truth depth, as well as camera intrinsics and extrinsics, and encompass a wide variety of dynamic scenes across both indoor and outdoor environments. Among them, PointOdyssey and DynamicReplica have additional 2D trajectory annotations for dynamic objects which can be used to construct pointmap matching ground truth following Eq. 2. All datasets are used to supervise geometry learning on Head1 and Head2, while only PointOdyssey, DynamicReplica, and TartanAir are used to train the proposed pointmap matching head.
Training and inference details. Our model architecture is based on the publicly available DUSt3R [52] model, utilizing the same backbone consisting of a ViT-Large encoder and a ViT-Base decoder. To fully leverage MonST3R’s geometry estimation capabilities in dynamic scenes, we initialize our model using the publicly available MonST3R checkpoint. For the newly introduced pointmap matching head, we initialize its weights from the pretrained Head2 weights of MonST3R. The temporal motion module is initialized following [14]. We train our network for 10 epochs with a cosine learning rate schedule, with an initial learning rate of 1e-4. In the first stage, which involves pairwise training, we use a batch size of 16 on 4 A100 GPUs (40G). In the second stage, where the temporal motion module is introduced, the batch size is set to 4 with a fixed temporal window length of 12. During each training iteration, we randomly sample a downstream task—3D point tracking, video depth estimation, or 3D reconstruction—to construct the input pairs and apply the corresponding loss function.
@@ 4.2 Video Depth Estimation
Following MonST3R [52] and CUT3R [43], we rescale all predictions from the same video to align them together by conducting two forms of alignment: per-sequence scale and shift alignment and per-sequence scale alignment. Thus, we can measure the per-frame depth quality and inter-frame depth consistency. We employ our proposed motion module for video depth estimation in a feed-forward manner as described in Sec.3.5 and compare our method against several variants of DUSt3R, including DUSt3R [44], MAST3R [26], MonST3R [52], Spann3R [41], and CUT3R [43]. Given 6 frames of 288×512 on an NVIDIA 4070 GPU, POMATO reconstructs the 3D point cloud in 0.7 seconds, whereas global alignment-based methods such as MonST3R require 5.8 seconds. As shown in Tab. 1, our method demonstrates comparable performance to the global alignment (GA)-based MonST3R [52] on the Sintel [4] and BONN [30] datasets, while surpassing it on KITTI dataset. Besides, we consistently outperform the state-of-the-art online method, CUT3R [43], across various settings. These results underscore the effectiveness of our approach, specifically (1) the joint learning of geometry and pointmap matching, and (2) the temporal motion module.
@@
Method L-12 L-24 L-12 L-24 L-12 L-24 L-12 L-24
SpatialTracker [47] 20.46 20.71 21.64 20.67 30.41 25.87 24.17 22.42
DUSt3R [44] 19.03 19.03 29.02 25.55 9.72 6.50 19.26 17.03
MASt3R [26] 16.58 17.35 27.36 26.46 11.78 8.09 18.57 17.30
MonST3R [52] 27.31 27.92 28.30 26.13 16.50 11.06 24.03 21.70
POMATO 33.20 33.58 31.57 28.22 24.59 19.79 29.79 27.20
Table 2: 3D tracking evaluation. We report the APD metric to evaluate 3D point tracking on the PointOdyssey [54], ADT [31], and PStudio [20] datasets. L-12 and L-24 indicate tracking within the temporal length of 12 frames and 24 frames, respectively.
@@
For 3D point tracking task, we use the Aria Digital Twin (ADT) [31], and Panoptic Studio (PStudio) [20] benchmarks from the TAPVid-3D [25] dataset along with the validation set on the PointOdyssey [54] dataset. We report the Average Percent Deviation (APD) metric, which quantifies the average percentage of points within a threshold relative to the ground truth depth. The APD metric serves as a direct measure of the accuracy of the predicted tracking. We reformulate the datasets and project all the query points within a temporal window to the first frame. We report tracking results on the length of 12 and 24 frames. As shown in Tab.2, our POMATO achieves the best performance on both PointOdyssey and ADT datasets. It’s worth mentioning that SpatialTracker [47] is a state-of-the-art network tailored for 3D point tracking with ground truth camera intrinsic as additional input data. POMATO surpasses it on two datasets and improves the average APD metric by 23.3% and 21.4% for 12 frames and 24 frames, respectively. For DUSt3R-based methods, we use the output of Head2 as tracking results. Obviously, the ambiguous matching representation limits its capability to handle this fine-grained 3D reconstruction task in dynamic scenes.
Video Depth Tracking (12 Frames)
Temporal Length Sintel [4] Bonn [30] KITTI [13] PointOdyssey [54] ADT [31] PStudio [20]
@@
Pair-wise 0.548 46.2 0.087 94.0 0.113 89.5 32.06 29.87 23.10
6 frames 0.436 51.3 0.076 95.9 0.085 93.5 32.69 30.93 24.52
12 frames 0.416 53.6 0.075 96.1 0.086 93.3 33.20 31.57 24.59
Table 3: Ablation study on the temporal motion module. The introduction of the temporal motion module brings a significant improvement. As the temporal window length enlarges from 6 frames to 12 frames, we obtain an overall consistent improvement.
@@
POMATO 0.020 0.010 0.509 0.037 0.016 1.782
Table 4: Pose estimation. Our method achieves an overall best performance and improves the RPE rot metric significantly.
Figure 7: Effectiveness of our pointmap matching head. Without explicitly filtering out the motion area, both pose and geometry estimation will be degraded.
@@ 4.4 Camera Pose Estimation
Following DUSt3R-based methods, we perform global alignment with the model trained in the first stage on the Bonn [30] and TUM [37] datasets. The sampling stride is set to 5 for the Bonn dataset and 3 for the TUM dataset. Compared with optical-flow assisted global alignment in MonST3R, the dynamic mask is computed according to Eq. 5 while the 2D pseudo label is replaced by projecting the pointmap matching results to 2D coordinates with estimated camera intrinsic. Absolute Translation Error (ATE), Relative Translation Error (RPE trans), and Relative Rotation Error (RPE rot) are reported. The evaluation results over 40 frames are reported in Tab. 4. Notably, POMATO obtains an overall state-of-the-art performance and significantly improves the RPE-rot metric, surpassing MonST3R by 55.4% and 13.3% on the TUM and Bonn datasets.
@@ 4.5 Ablation Study
Bonn [30] PointOdyssey [54] ADT [31] PStudio [20]
Method ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$ APD $\uparrow$ APD $\uparrow$ APD $\uparrow$
W/O Head3 0.040 0.015 1.721 29.10 29.62 16.94
W/ Head3 0.037 0.016 1.782 32.06 29.87 23.10
Table 5: Ablation study on the effectiveness of the pointmap matching head. The comparisons are reported on the pose estimation and 3D point tracking tasks.
We conduct extensive ablation studies to evaluate the effectiveness of the temporal motion module and the proposed pointmap matching head. As shown in Table 3, we report results for three models: one trained with only pairwise images (first-stage training), one using a shorter temporal window of 6 frames, and another using the default temporal window length of 12 frames. Incorporating temporal consistency yields substantial improvements across all datasets for video depth estimation and 3D point tracking. Further improvement is achieved when the temporal window length increases from 6 frames to 12 frames. In Table 5, we evaluate the effectiveness of the pointmap matching head. While it introduces only a modest improvement in the ATE metric, we attribute this to the limited motion and minimal viewpoint variation in the indoor evaluation dataset. As illustrated in Fig. 7, under challenging in-the-wild conditions with significant motion and rapid viewpoint changes, removing the pointmap matching head introduces ambiguity in explicit rigid transformation estimation, resulting in a clear degradation in performance. To further demonstrate the impact of the pointmap matching head on 3D point tracking, we conduct tracking experiments over 12 frames using the pairwise input setup. Clearly, removing the pointmap matching head (using only Head2) leads to an inevitable performance drop, emphasizing explicit correspondence modeling for reliable long-term tracking.
@@ 5 Discussion and Conclusion
We introduce POMATO, a unified framework for geometry estimation and motion understanding in dynamic scenes. By leveraging the proposed pointmap matching head, our method effectively distinguishes moving regions, thereby mitigating the interference introduced by dynamic objects. The temporal motion module further facilitates the learning of temporal dynamics across frames, enhancing scale consistency and improving performance in tasks where both geometry and matching are critical—most notably, 3D point tracking. The downstream temporal tasks including 3D point tracking, video depth estimation, and 3D reconstruction can be easily applied in a feed-forward manner. In future work, we plan to scale up training with more dynamic reconstruction and matching datasets to further enhance 3D reconstruction and tracking performance.
Acknowledgement. This work was supported by the National Natural Science Foundation of China (No. 62206244)
@@
[…]